Which of the ordered pairs (3,1), (0,-4), (-3,-7) are solutions for the equation x-y=4?
Please help I dont understand how to resove this.
Thank you

Answer by ankor@dixie-net.com(22740)

Which of the ordered pairs (3,1), (0,-4), (-3,-7) are solutions for the equation x - y = 4?
:
Since they give the possible answers it's very easy. The ordered pairs represent x and y:
:
Substitute the given values for x and y and see if the equation is true
Take the 1st ordered pair: x = 3; y = 1
x - y = 4
3 - 1 = does not equal 4 so this is not the solution:
:
2nd ordered pair x = 0, y = -4
0 - (-4) = 4; a minus a minus is a plus, remember
0 + 4 = 4 so this is a solution
:
3rd ordered pair: x = -3, y = -7
-3 - (-7) = 4
-3 + 7 = 4 is true so this is a solution also
:
Did this help you?